OpenStudy (anonymous):

Help please? Will give medal. The following table shows the values of f(x) and g(x) for different values of x: x 1 2 3 4 5 f(x) 3 9 27 81 243 g(x) 4 4 4 4 4 Which of the following best describes the function f(x) + g(x)? Choices: 3x + 4 3x + 4 3x + 4x 4(3x)

jimthompson5910 (jim_thompson5910):

Notice how f(x) is the list of powers of 3 3 = 3^1 9 = 3^2 27 = 3^3 81 = 3^4 243 = 3^5 etc etc

jimthompson5910 (jim_thompson5910):

g(x) is always 4 no matter what x is

jimthompson5910 (jim_thompson5910):

which means \[\Large f(x) + g(x) = 3^{x} + 4\]

OpenStudy (anonymous):

its the third choice right?

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

i just looked at it, thanks c: